OpenStudy (anonymous):

how do i solve by factoring 2a^2+3=79? i dont know what steps i need to take or how to show my work on this problem.

OpenStudy (agreene):

\[2a^2+3=79\] subtract 79 from both sides: \[2a^2-76=0\] factor \[2(a^2-38)=0\] divide by 2 \[a^2-38=0\] add 38 to both sides \[a^2=38\] square root \[a=\pm\sqrt{38}\] So: \[\large a=\sqrt{38} ;-\sqrt{38}\]